OpenStudy (k.binks):

Write the equation of a circle with a center at (-6, 9) and a radius of 12?

Vocaloid (vocaloid):

let (-6,9) be (h,k), and let the radius = r use the formula (x-h)^2 + (y-k)^2 = r^2

OpenStudy (k.binks):

So, (x+6)^2 + (y-9)^2=12^2? Do I change the 12^2 to 144?

OpenStudy (solomonzelman):

yes, you change 12 to 144, and then your complete equation is?

OpenStudy (solomonzelman):

i mean 12² to 144

OpenStudy (solomonzelman):

u get?

OpenStudy (k.binks):

I get. (x+6)^2 + (y-9)^2= 144, right?

OpenStudy (solomonzelman):

yes, that is right

OpenStudy (solomonzelman):

and that is it as well. you are done

OpenStudy (k.binks):

Thank you!

OpenStudy (solomonzelman):

anytime!